OpenStudy (anonymous):

S=2πr^2+2πrh solve for π

OpenStudy (mathlegend):

Can't you factor out π?

OpenStudy (anonymous):

I am totally clueless on this. Terrible at math, taking a developmental course to improve, but the instructor likes to give work he has not covered to see what we can do.

OpenStudy (mathlegend):

Well, I would consider getting up to par with factoring especially if he is giving stuff like this. You can factor out the π since it is in both terms. S=2πr^2+2πrh S = π(2r^2+2rh) Does that make sense so far? We took out π...

OpenStudy (anonymous):

I am with you so far. Makes sense to see it, you make it look easy.

OpenStudy (mathlegend):

S = π(2r^2+2rh) When a number or variable is touching the parentheses this means multiply... so now we must divide to get pi by itself... \[\frac{ S }{ (2r^2+2rh) }=\frac{ \pi(2r^2+2rh) }{ (2r^2+2rh) }\]

OpenStudy (mathlegend):

Now, you see that they will reduce on the right side... which gives us...

OpenStudy (mathlegend):

\[\frac{ S }{ (2r^2+2rh) }= \pi\]

OpenStudy (mathlegend):

Make sense?

OpenStudy (anonymous):

I am shaking my head at the simplicity of the explanation. You are truly a math legend. How many thank you's can I send your way?

OpenStudy (mathlegend):

:) Glad I could help! Welcome to OpenStudy!!!

OpenStudy (anonymous):

Thank You! You have helped more than you know!